If 465cm³ of SO2 can diffused through a porous partition in 30secs , how long will 620cm³ of hydrogen sulphide take to diffuse through the same (H=1,S=32,O=16)


1
Expert's answer
2022-06-02T16:54:04-0400

Solution:

Graham's law states that the rate of diffusion or of effusion of a gas is inversely proportional to the square root of its molecular weight.


Graham's Law can be expressed as:

rate ∝ 1 / (molecular weight)1/2

Therefore,

(r1 / r2) = (M2 / M1)1/2


r1 = rate diffusion of SO2 = (465 cm3) / (30 s) = 15.5 cm3 s−1

r2 = rate diffusion of hydrogen sulphide (H2S) = 620 cm3 / t

M1 = molecular weight of SO2 = 32 + 2×16 = 64 g mol−1

M2 = molecular weight of H2S = 2×1 + 32 = 34 g mol−1


Thus:

(15.5 / r2) = (34 / 64)1/2

15.5 / r2 = 0.729

r2 = 21.26 cm3 s−1


r2 = 21.26 cm3 s−1 = 620 cm3 / t

t = 29.16 s


Answer: 29.16 seconds